Would someone please explain the steps to solve for y?
y-1/x=3/5
Thank you

Answer by checkley77(12844)

y-(1/x)=3/5 add 1/x to both sides.
y=3/5+1/x
OR:
(y-1)/x=3/5 cross multiply
5(y-1)=3x
5y-5=3x add 5 to both sides.
5y=3x+5 divide both sides by 5.
y=(3x+5)/5 ans.